OpenStudy (anonymous):

how do you integrate x^2cos(x^3)dx

OpenStudy (therealmeeeee):

Let z=x² so dz=2xdx The integral becomes: ∫1/2 z cosz dz Using by parts u=z, dv=cosz du=1,v=sinz So ∫1/2 z cosz dz =1/2 [zsinz - ∫sinzdz] =1/2 [zsinz+cosz] + c =1/2(x²sinx² + cosx²} + c

OpenStudy (therealmeeeee):

do this help

OpenStudy (anonymous):

yes im honestly just stuck on what to do with cos(x^3) because i know what to do with the other stuff (meaning integration by parts)

OpenStudy (anonymous):

and im kinda lost on what z and dz is because i use u,du dv,v

OpenStudy (therealmeeeee):

yes

OpenStudy (therealmeeeee):

dont use that one no more

OpenStudy (isaiah.feynman):

I hope you didn't copy this "solution" from somewhere online! @TheRealMeeeee

OpenStudy (therealmeeeee):

no i didnt not this time

OpenStudy (isaiah.feynman):

Hard to believe. You typed so much in a few moments... can you even EXPLAIN your solution?

OpenStudy (therealmeeeee):

well my computer is mesed up it didnt show how long i was typing till the end

OpenStudy (isaiah.feynman):

Yeah right.

OpenStudy (therealmeeeee):

ok watever

OpenStudy (anonymous):

my real question is how do you integrate cos(x^3) because i know that is what you use for dv

OpenStudy (anonymous):

and i know you got that from yahoo i saw it before

OpenStudy (therealmeeeee):

ok watever

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how do i integrate cos(x^3)

OpenStudy (anonymous):

because i know its by parts

OpenStudy (perl):

just set u = x^3 , then du = 3x^2

OpenStudy (anonymous):

instead do this: u = x^3 du = 3x^2 dx \[\int\limits_{}^{}x^2 \cos(x^3) dx = \int\limits_{}^{}\frac{ 1 }{ 3 }\cos(u)du = \frac{ sinu }{ 3 } + C = \frac{ \sin(x^3) }{ 3 } + C\]

OpenStudy (anonymous):

thanks you!

OpenStudy (perl):

@TheRealMeeeee youre answer is wrong. read the problem carefully

OpenStudy (isaiah.feynman):

See the way its done? @TheRealMeeeee

OpenStudy (anonymous):

Euler has the correct answer, a simple sub is all you need.

OpenStudy (therealmeeeee):

yes i see

OpenStudy (anonymous):

thank you i thought i had to do integration by parts